### v1

# Continuous-Time State Estimation Methods in Robotics: A Survey

Abstract
Accurate, efficient, and robust state estimation is more important than ever in robotics as the variety of platforms and complexity of tasks continue to grow. Historically, discrete-time filters and smoothers have been the dominant approach, in which the estimated variables are states at discrete sample times. The paradigm of continuous-time state estimation proposes an alternative strategy by estimating variables that express the state as a continuous function of time, which can be evaluated at any query time. Not only can this benefit downstream tasks such as planning and control, but it also significantly increases estimator performance and flexibility, as well as reduces sensor preprocessing and interfacing complexity. Despite this, continuous-time methods remain underutilized, potentially because they are less well-known within robotics. To remedy this, this work presents a unifying formulation of these methods and the most exhaustive literature review to date, systematically categorizing prior work by methodology, application, state variables, historical context, and theoretical contribution to the field. By surveying splines and Gaussian processes together and contextualizing works from other research domains, this work identifies and analyzes open problems in continuous-time state estimation and suggests new research directions.
Index Terms:
State Estimation, Continuous-Time, Splines, Gaussian Processes, Optimization, Sensor FusionI Introduction
Autonomous driving, extra-planetary exploration, infrastructure inspection, and environmental monitoring are examples of the complex tasks expected of robotic platforms today. To cope with these diverse challenges, onboard sensors have increased in quantity, variety, and bandwidth, and the requirement for accurate and robust state estimation is more critical than ever. The constraints of these platforms, including cost, size, sensor quality, power consumption, and computational resources, vary significantly. Some platforms, like commercial drones or handheld devices, are highly constrained, while autonomous cars, construction machines, and others may have access to more onboard compute. As the demand for such diverse robotic systems grows in the coming decades, so will the requirement for general, flexible, and scalable state estimation solutions. A paradigm shift, from discrete to continuous-time, may be the key to achieving this.
Discrete-Time (DT)
Historically, state estimation techniques have operated in discrete-time, which means the state of the robot or process is estimated at specific moments in time. However, the evolution of the system is not modeled, and intermediate states cannot be inferred. These times must include the measurement times, so regardless of how quickly the system state evolves, DT methods require estimating variables at all of these sample times [1, 2, 3]. The computation of such algorithms scales poorly as the number and frequency of sensors increase, compromising their ability to run in real-time. Several techniques and tricks are employed to work around this issue. One of the most common is to aggregate measurements over time, especially from sensors that capture hundreds or thousands of measurements per second, to form lower-frequency pseudo-measurements. Examples include the accumulation of points from a sweeping Light Detection And Ranging (LiDAR) sensor for inter-scan or scan-to-map registration [4, 5, 6, 7, 8], feature extraction in event [9] or Rolling Shutter (RS) cameras, and Inertial Measurement Unit (IMU) preintegration [10, 11, 12, 13, 14]. Motion distortion may occur if the platform moves substantially over this accumulation period. Motion Distortion Correction (MDC) methods such as LiDAR deskewing [15] or RS compensation [16] attempt to correct for this. However, MDC is fundamentally a chicken-and-egg problem since the motion-corrected measurements are required to estimate the motion accurately. Consequentially, these methods introduce hard-to-model errors into the system. Some DT systems (e.g., [17]) also achieve state reduction by triggering sensors to capture simultaneously and often have complicated hardware triggering systems and clock synchronization in place to ensure that the timestamps reported by sensors are well-aligned to a central clock. Yet, in systems without proper time synchronization, estimating time offsets between sensor clocks is a complex task for DT state estimators [18, 19].
Continuous-Time (CT)
Continuous-time methods (Figure 1) differ by modeling the underlying process as a continuous function of time and have the defining characteristic that the state can be queried at any time (sometimes within certain bounds). This capability can, on the one hand, be useful for downstream tasks that rely on accurate state estimates, such as planning or control. On the other hand, it allows for querying interpolated states at measurement times during estimation without requiring additional explicitly defined variables. By providing a layer of abstraction between the time-varying state and the optimization variables, CT methods allow the number of variables to be flexible and their generation to be application and sensor-agnostic. This property facilitates an elegant means of trading off accuracy and computation time while considering the robot’s dynamics. Such a system encourages the direct inclusion of high-frequency measurements, such as from an IMU or LiDAR, without necessarily creating new estimation variables, avoiding the errors introduced by measurement aggregation. Furthermore, by design, these systems can handle asynchronous sensor data and can naturally jointly estimate the time offset between sensor clocks by shifting the interpolation time [19].
Significance of this Work
The CT paradigm could be the key to accurate, robust, and scalable state estimation. Indeed, CT methods are already being leveraged to achieve state-of-the-art performance across a full spectrum of applications. However, they remain relatively niche and less well-understood within the broader robotics community. The field has progressed substantially since the last CT survey almost a decade ago [20]. Other review works have focused more specifically on a single method [19] or sensor modality [21], or on comparative performance [22]. To update the community on modern CT methods and lower the bar of entry to future research, this work makes following contributions:
A concise background (Section II) and consolidated formulations for the established CT methods (Section III).
The most complete survey of CT state estimation to date (until October 2024), categorized by method, application, state variables, historical context, or theoretical contribution, as most appropriate (Section IV).
Identification and analysis of open problems (Section V).
A discussion of the scope of CT methods in other robotics’ domains and applications (Section VI).
II State Estimation in Robotics
II-A Preliminaries
This section provides a brief introduction to the field of state estimation and establishes a common notation for DT estimation and the CT methods introduced in Section III. In the interests of conciseness, a detailed theory of optimization-based state estimation theory will not be covered as has been done in prior works [23, 24, 25, 26]. This work considers state estimation problems which aim to tractably estimate the probability distribution of a (finite) set of (continuous) random variables that parameterize a system of interest. Lie theory [27, 28, 29] has proven to be a valuable framework for treating these variables accurately in the context of state estimation in robotics; in particular, the and matrix Lie groups for representing -dimensional orientations and poses, and for monocular camera poses where scene scale is unobservable. Because Lie groups are smooth, differentiable manifolds, they have unique tangent spaces, which are vector spaces amenable to the linear algebra operations required during optimization. This work will adopt the notation and right-handed conventions of Sola et al. [29].
States
A continuously evolving process is defined by two components; i) a time-varying state, denoted as for time variable , such as the pose of a robot, and ii) its time-invariant context, a set of variables denoted as , such as sensor calibration or map parameters. Without yet proposing a relationship to , a set of variables is defined to describe . Estimating these variables, collected together as , will then be possible through measurements and prior knowledge.
Measurements
A measurement model describes how an observation is generated from one or more sensors. The measurement function for this model, , defines a (measured) observation as a function of the state and associated noise , usually assumed to be drawn from a Gaussian distribution and independent between measurements. Common examples in robotics include range-bearing, projective geometry, kinematic, and inertial measurement models. A residual function is defined to obtain a vector-valued residual , which is commonly subtraction or the generalized operator [29, Equation (26)] for Lie groups. Alternatively, and may be implicitly combined such as in point-to-plane [5, 30] or on-unit-sphere angular [31] distances.
Prior Knowledge
Prior models incorporate assumptions about a system before evidence from measurement models is incorporated. They can be formulated in the same way as measurement models, except that is not observed but set a priori. In practice, a belief about the system’s initial conditions (e.g., pose and velocity) is often required to ensure a unique solution in the optimization. For a more complex example, planar surface regularization may be applied if an environment is known to be highly structured. As with measurement models, residuals can be defined for such priors.
Process Models
Process models (a.k.a. system dynamics, motion models) describe the evolution of the state over time. Process models have frequently appeared in DT filters, occasionally in optimization-based approaches, and are a fundamental idea used in some CT methods. Since the process model is known a priori and does not require measurements, it can be seen as a type of Bayesian prior. If the process is Markovian, then it can be formulated in terms of the current process state , process inputs , zero-mean process noise , and context , as a Nonlinear Time-Varying (NTV) Stochastic Differential Equation (SDE)
Commonly, does not depend on the time-invariant parameters and is linearized about a current estimate of the other inputs:
The general solution of this Linear Time-Varying (LTV) SDE [32, 33] (Itô calculus not required here due to linearity), given some initial state at time , is
Here, is the transition matrix for the process, with properties for any , and . It can be found analytically for simple LTV-SDEs but can be challenging to find in general, and so without an analytical solution, it must be computed numerically [34, 35, 36].
Maximum A Posteriori (MAP) Estimation as Least-Squares Optimization
Modern state estimation approaches are fundamentally probabilistic, relying on the principle of Bayesian inference. The application of Bayesian theory to DT state estimation in robotics has been covered thoroughly in prior work [25, 26, 37, 38], and the findings transfer to CT state estimation as well. One powerful consequence of this theory is that under the assumption that all residual errors (e.g., from measurement, prior, and process models) are drawn from zero-mean Gaussian distributions, an optimal estimate (in the MAP sense) of the state is obtainable by minimizing the sum of their squared Mahalanobis distances. Specifically, for residuals with individual covariances , a Gaussian approximation for the posterior distribution (a.k.a. Laplace approximation) is found [39] by solving
II-B Discrete-Time (DT) State Estimation
In Discrete-Time (DT) state estimation, the process state variables to be estimated are defined as the process state itself at a specific set of times , i.e., . There are two broad categories of approaches for solving DT state estimation problems: filtering and smoothing.
Filtering
Bayesian filtering adopts a predict & update pattern that relies on the Markov assumption. This assumption states that the evolution of a process, such as robot motion, can be inferred from its current state only and does not depend on how this state was reached. In the predict step, the prior belief of the state is propagated according to a process model. Then, in the update step, measurements are used to estimate the posterior distribution for the state. Kalman Filters (KFs) [40, 41, 42] and Information Filters (IFs) [37] have proven to be highly effective and efficient, and are still widely used, especially when latency and computational load must be low. If the state has a non-Gaussian distribution, discretized state-space methods like the histogram filter, or Monte Carlo methods such as the Particle Filter (PF), which uses samples to represent the state probability distribution, are often used instead [37].
Smoothing
Bayesian smoothing estimates the joint probability distribution of variables representing (a window of) the current and previous robot states, typically relying on optimization techniques. Smoothing methods have gained popularity in many robotics problems as they have matured in recent decades regarding robustness, accuracy, and speed. Due to the Markov property, variables involved in the optimization are usually related to only a small subset of (neighboring) variables. These approaches demonstrate speeds compatible with onboard robot operation by exploiting this inherent (near diagonal) sparsity through efficient linear algebra tools (e.g., sparse LU, LDL, Cholesky, and QR factorization), which are the basis of efficient optimal solutions of linear problems and fast iterative solvers based on relinearization for nonlinear problems (e.g., Gauss-Newton (GN) [24], Levenberg-Marquardt, Steepest Descent, Powell’s Dogleg) [25].
II-C Continuous-Time (CT) State Estimation
Since processes in the physical world usually change continuously over time, it is natural to seek a continuous-time representation to model them. The fundamental property of CT methods is their ability to query the state at any required time , given a set of variables . The exact relationship between and depends on the chosen method. Yet, a key advantage of these methods is that the number and temporal placement of variables are flexible. Intuitively, the quantity of variables must scale with the duration of the process and needs to be increased to accurately model more complex dynamics. Notably, these methods do not require variables for every measurement time and can naturally accommodate asynchronous sensor data through their inherent ability to interpolate (or extrapolate). To remain computationally competitive with DT methods, CT methods rely on the local support property, meaning that inferring a state at a particular time depends only on a small number of (local) variables.
The robotic research community has experimented with several CT methods for state estimation over the past years. i) Interpolation (e.g., Linear Interpolation (LI)) and integration methods (Section III-A) appear frequently, in part due to their simplicity and low computational complexity. ii) Temporal splines (Section III-B), especially B-splines, are among the most popular CT methods to emerge in robotics. Through control points and knots, splines embrace the abstraction of state inference from estimated variables. iii) ‘Exactly Sparse’ Temporal Gaussian Processes (TGPs)111 These are not to be confused with sparse GPs [43], which address the cubic complexity of GPs by selecting active points as a subset of training points or optimizable pseudo-inputs that account for the entire training set [44]. (Section III-C) are a recent, compelling approach for CT state estimation. Table I provides a high-level overview.
| Specifications | Linear Interpolation | Temporal Splines | Temporal Gaussian Processes |
| Interpolation | linear | polynomial | process model |
| Lie Group Support | ✓ | ✓ | ✓ |
| Covariance Interpolation | ✓ | Section V-A5 | ✓ |
| Extrapolation | ✓ | [45] | ✓ |
| Non-Interpolated Residual Support | ✓ | ✗ | ✓ |
| Initialization | linear extrapolation | Section V-A4 | process model |
| Process Variables | explicit states | control points | explicit (Markovian) states |
[…]
| Variable Size | size of | size of | (size of ) + (DoF of ) (derivatives) |
| Variables per Interpolation | 2 | 2 | |
[…]
III Theory
III-A Interpolation & Integration
When interpolation and integration are discussed in the context of CT state estimation, the state variables precisely represent the process state (i.e., ). Interpolation is defined as the inference of intermediate states from estimated curves that pass exactly through these explicit states.222 However, the term interpolation is frequently overloaded, and so is used more liberally in later sections to describe inference with regression models that do not pass through every data point. This is common in the spline literature, although only some spline types are interpolating. On the other hand, integration uses rate measurements, such as the angular rate from an IMU, for this inference.
III-A1 Linear Interpolation (LI)
One of the simplest and fastest interpolation methods is Linear Interpolation (LI). The generalized LI of two adjacent states can be written as
where are the addition and subtraction operators for composite manifolds [29, Section IV], and for interpolation time (outside is extrapolation). For Lie groups, these operators become [29, Equations (25) and (26)], which when applied to quaternions [46] or rotation matrices [24] yield spherical linear interpolation (SLERP). This has also been applied to poses [24]. When the states belong to vector space, this simplifies to
Importantly, LI maintains a constant rate of change (‘velocity’) while other interpolation methods, such as Quaternion Linear Blending (QLB) [46] do not. LI therefore enables rapid inference of the state at any time under this constant-rate assumption, and allows the number of optimized states to be chosen flexibly. While nonlinear interpolation methods exist, they are usually captured within the scope of Temporal Basis Functions (TBFs), the foundation for temporal splines [20] and TGPs [47]. Indeed, LI is exactly a special case of these methods. Despite the development of more advanced methods, LI remains popular for its simplicity and speed, and may be sufficiently accurate to model many processes.
III-A2 Numerical Integration
Numerical integration in robotics is sometimes discussed in the CT state estimation literature, as it allows inference of intermediate states from rate measurements. However, unlike the other approaches, numerical integration does not provide a closed-form state inference function. In robotics, this is often used for the motion compensation of scanning sensors such as LiDARs or RS cameras, or for IMU preintegration. A common method is Riemann summation (a.k.a. the rectangle rule), which assumes a constant rate of change and hence is closely related to LI. When formulated as an Ordinary Differential Equation (ODE), numerically integrating constant rates over short time intervals is known as the forward Euler method. Regressing and integrating Gaussian Processes for IMU preintegration has also been explored (Section IV-A6).
III-B Temporal Splines
A number of different formulations for splines have been presented in the literature [31, 46, 20, 48, 49, 50, 51, 52, 53, 45]. The formulation presented in this work is primarily based on the works of Furgale et al. [20] and Sommer et al. [53]. A spline is a function of a scalar input whose resultant curve (in vector or manifold space) is a piecewise polynomial of degree . The generated curve is composed of segments. The order is a fundamental property of a spline guaranteeing certain continuity properties and is related to the degree as . In temporal splines, the input scalar values represent time.333 Technically, splines are a special case of Temporal Basis Functions where the bases are the polynomials of degree 0 to .
A set of control points (the estimation variables) and a set of scalar times called knots characterize a temporal spline. Each spline segment depends only on control points. This is the local support property for splines and is key for efficient computation and optimization. The -th polynomial segment (indexing from ) depends on a subset of the control points and is defined on the interval .444 To capture the degenerate case in a manner consistent with higher orders, define , and when no knots are defined, . A time lying within this interval can be normalized as
Polynomial powers of this normalized time are then stacked into the vector .
III-B1 Interpolation
Vector Space
Through , the blending matrix, one can compute the coefficients of the blending vector , that enable interpolation on the -th segment of the curve. Interpolation on the -th segment is simply
Figure 2 illustrates some 2-dimensional vector-space splines.
Lie Groups
To efficiently work with Lie group splines, the cumulative interpolation formulation must be used [53]. The cumulative blending matrix is defined as a cumulative sum over columns of , , and are the coefficients of the cumulative blending vector . Then, using the composition operator (matrix multiplication for many groups), the Exp map, and the Log map defined by [29, Equations (1), (21) and (22)], the interpolation at is given as
III-B2 Spline Types
The computation of the as-yet-undefined blending matrix depends on the spline type, defined by a selection of internal constraints that determine its continuity properties. The temporal continuity of a spline refers to its degree-of-differentiability, where is the number of derivatives that are continuous over the entire input domain of the spline. The -th derivative is a piecewise polynomial of degree . Closed-form formulas have been derived for the first three temporal derivatives [53]. also depends on the knots, which often have uniform spacing at some knot interval. Otherwise, the spline is non-uniform.
B-Splines
The most popular spline used in robotics applications is the B-spline, since it has the highest continuity possible; (not , e.g. in [19, 53, 54]). If the B-spline is uniform, then each element at indices of the blending matrix can be set using binomial coefficients [53] as
The blending matrix of non-uniform splines depends on knots (not , e.g., in [20]). It is the transpose of the basis matrix from Haarbach et al. [46], originally Qin [52], and so it can be computed by transposing the result of Qin’s recursive formula. Table II summarizes common spline types.
| Spline | Continuity | Interpolating | Example Applications | Additional Notes |
| Bézier | / | Every -th | Vector Graphics, Fonts | if control points are mirrored about knot points. |
| Hermite | / | All | Animation | Velocity set at control points. Special case of Bézier spline. |
| Kochanek-Bartels | / | All | Computer Graphics | Modification of Hermite Spline. |
| Cardinal | All | Animation | Special case of Kochanek-Bartels spline. | |
| Catmull-Rom | All | Animation, Smoothing | Special case of Cardinal spline. | |
| Linear | All | Fast Interpolation | Special case of spline with order . | |
| B-Spline | None in general | Animation, Camera Paths | Minimal support w.r.t. order, smoothness and domain. | |
| NURBS | None in general | Computer-Aided Design (CAD) | Special case of B-spline. |
III-B3 Optimization
In formulating a spline optimization problem, all priors and measurements must be included as interpolated factors. This means that interpolation (Equation 11) must be used in the residual computation to obtain the state or its temporal derivatives at the required sample times. Therefore, sufficient control points and knots are required to perform interpolation at all prior and measurement times. As a result, these factors in the optimization are associated with i) the control points of the spline segment(s) upon which interpolation will be performed, and ii) any necessary time-invariant variables. Jacobians for these factors, required for optimization, can be computed analytically [53, 55, 56, 22].
III-C Temporal Gaussian Processes (TGPs)
In contrast to splines, which model the state parametrically as a weighted combination of control points, TGPs are non-parametric, using instead a Gaussian Process (GP)
Here is the prior mean function and is the prior covariance function, or kernel. Interestingly, TGPs can be considered a weighted combination of infinitely many Temporal Basis Functions (TBFs), made non-parametric by substituting the mean and covariance with functions (known as the kernel trick) [57, 47]. Indeed, despite different design choices, CT methods (based on TBFs) are sometimes related. For example, posterior interpolation with a WNOA prior (see Section III-C2) yields a cubic Hermite spline interpolation scheme [24]. As in DT state estimation, the estimated process variables represent the state at specific times. Similarly to spline knots, the times of these support states may be chosen freely (uniformly or non-uniformly).
This method requires the evolution between these support states to be modeled as a LTV-SDE process as introduced in Section II. While this imposes a Markovian restriction on the CT state, augmenting the base state with derivatives to overcome this is usually straightforward, shown later in Section III-C2. Since many systems of interest in robotics can be modeled as LTV-SDEs, this method is quite general.
It is assumed that the process is governed by CT white noise
where is the (symmetric, positive-definite) power-spectral density matrix and is the Dirac delta function. The mean function can be derived from Equation 5 as
From this it is easy to formulate a residual between two explicitly estimated states at times and as
The covariance function can also be derived from Equation 5 (cf. [34, 35, 24, 58]) as
Together, the residual with corresponding prior covariance are used as a binary factor between adjacent states, commonly referred to as the GP (motion) prior factor.
III-C1 Interpolation
Vector Space
The key insight is that the process model has been used to define the prior mean and covariance functions of the GP. Because the process model is Markovian, the inverse covariance is exactly sparse (block-tridiagonal), and by exploiting this sparsity [34, 35], the interpolated posterior mean and covariance can be derived555 The GP interpolation equations can also be derived from the perspective of optimal control theory [22, appendix A]. for as
Importantly, interpolation relies only on the two temporally adjacent states. If as in commonly used models, then Equation 22 further simplifies to
Lie Groups
To work with Lie group states, which would yield a nonlinear SDE process in almost all non-trivial cases, the method of local LTV-SDEs is used [59, 60, 61], which utilizes the local tangent space of the manifold. Let be the global state, where is a Lie group element (e.g., pose), and are its first and second temporal derivatives. Between every pair of explicitly estimated states, and , the NTV-SDE is approximated as an LTV-SDE for through the use of a local mapping
where is the right Jacobian of the manifold as defined by [29, Equation (67)].666 Note that does not have a closed form in general, but can be approximated [62]. A closed form solution exists for [63]. The local-to-global mapping is
Therefore, one stores and optimizes the global state variables but uses the local LTV-SDE for interpolation. Interpolating the state mean at is then a three-step process:
1.
Map explicitly estimated global states to local states using Equation 27.
2.
Interpolate local state using Equation 22 (or Equation 26).
3.
Map the interpolated local state to a global state using Equations 28 and 29.
A similar procedure is available for the covariance [36].
III-C2 GP Priors
White-Noise-On-Acceleration (WNOA)
The WNOA or ‘constant-velocity’ prior is a generic process model and was the first to be used for GP-based CT state estimation [57]. Other priors are discussed in Section IV-C. It assumes the second derivative (‘acceleration’) is driven by white noise (Equation 16). For vector-valued , this is written as
A first-order SDE (satisfying the Markovian condition) is formulated by including the first derivative in the state:
This yields a Linear Time-Invariant (LTI) SDE with
The transition, covariance and precision matrices for the LTI-SDE are calculated in closed form (with ) as
Since , the computation of the GP prior residual is simplified, and Equation 26 can be used.
If is a Lie group object, then one must have the global Markov state . The local Markov state defined at and its temporal derivative are
The local WNOA prior thus assumes that acceleration in the local frame is driven by white noise, as . Substituting into Equation 19, the WNOA prior residual (with ) can be written as
IV Survey
IV-A Interpolation & Integration
An overview of the considered works that use interpolation and integration in a CT formulation is provided in Table III.
IV-A1 Linear Interpolation (LI) in LiDAR and RADAR Systems
Among the earliest works that formulate LI as a CT method are the LiDAR SLAM system by Bosse and Zlot [64] and Zebedee [65], which apply the constant-velocity assumption over short time intervals for pose interpolation in their surfel-based LiDAR and LiDAR-inertial Simultaneous Localization And Mapping (SLAM) systems, respectively. Due to its speed, many works have adopted LI for point cloud Motion Distortion Correction (MDC) in LiDAR odometry [66, 67, 68, 4, 5, 69, 70], LiDAR-inertial odometry [71, 72, 73], LiDAR-inertial SLAM [74], LiDAR-visual-inertial odometry [75], and inertial-wheeled odometry [76]. Ceriani et al. [77] model an trajectory with LI in their LiDAR SLAM system, repeatedly correcting for motion distortion in its scan-to-map ICP-based optimizer. FAST-LIO [78] performs a Riemann-style summation of IMU measurements in its LiDAR-Inertial Odometry (LIO) system, both forward, for state propagation in the Kalman Filter (KF), and backward, for point cloud MDC. LI has also been applied to MDC of spinning RADAR [79, 80].
The LiDAR Odometry (LO) systems CT-ICP [81] and ECTLO [82] embrace the CT formulation for LiDAR MDC within the residuals of iterative optimization, as does the LiDAR-inertial SLAM system by Park et al. [83] in its surfel and inertial residuals. Later surfel-based SLAM works by the same authors [84, 85] maintain LI as the primary representation for fast inference, but use cubic B-splines to obtain pose corrections. Park et al. [86] apply LI over poses to the spatial-temporal calibration of LiDAR-camera systems. Recently, Zheng and Zhu [30] formulate point-to-plane residuals for every point of a scanning LiDAR in their LO system, using LI to estimate the pose at the acquisition time of each point.
IV-A2 Linear Interpolation (LI) in Camera Systems
Many works have also applied LI to camera systems, especially in Rolling Shutter Compensation (RSC). An early example was by Klein and Murray [87], who applied LI to perform RSC within PTAM [88]. The constant velocity was determined using the feature correspondences, and a similar idea was adopted by the RANSAC approach of Anderson and Barfoot [89]. RSC was very common in early CT works, through LI [90, 91, 92, 93, 94, 95] or Riemannian integration of gyroscope measurements [96]. LI has also been used to efficiently infer poses for multi-sensor fusion [97], as well as in multi-camera odometry and calibration systems [97, 98].
IV-A3 Arguments for Linear Interpolation (LI)
A common technique present in many of the CT LI works [64, 65, 99, 100, 101, 84, 85], is the separation of a base trajectory of discrete poses and a correction trajectory during estimation. The key idea is that when using LI, corrections can be adequately represented at a lower frequency than the poses themselves [101, 84], with few coefficients to be optimized. Furthermore, LI is simple to implement with low computational cost.
| Authors | Year | Applications | Time Invariant Estimates | Time-Varying Estimates | |||||||
| State(s) | State(s) | State Representation(s) | |||||||||
| Ait-Aider and Berry [90] | 2009 | RS Object Tracking | Linear Interpolation | ||||||||
| Bosse and Zlot [64] | 2009 | LiDAR SLAM | Linear Interpolation (& Cubic Spline) | ||||||||
| Klein and Murray [87] | 2009 | RS SLAM | Riemann Summation | ||||||||
| Forssén and Ringaby [91] | 2010 | RSC | - | Linear Interpolation | |||||||
| Hong et al. [66] | 2010 | PC MDC for PC Registration | - | Linear Interpolation | |||||||
| Moosmann and Stiller [67] | 2011 | PC MDC for LO | - | Linear Interpolation | |||||||
| Ringaby and Forssén [92] | 2011 | RSC | - | Linear Interpolation | |||||||
| Karpenko et al. [96] | 2011 | RSC for Video Stabilization | Linear Interpolation | ||||||||
| Hedborg et al. [94] | 2012 | RSC for SfM | Linear Interpolation | ||||||||
| Bosse et al. [65] | 2012 | LiDAR-Inertial SLAM | Linear Interpolation | ||||||||
| Jia and Evans [93] | 2012 | RSC for Video Rectification | - | Linear Interpolation | |||||||
| Anderson and Barfoot [89] | 2013 | LO | - |
| Vivet et al. [79] | 2013 |
| Dong and Barfoot [68] | 2013 | PC MDC for PC Registration for LO | - | Linear Interpolation | |||||||
| Li et al. [102] | 2013 | VIO | - |
| Guo et al. [95] | 2014 | RS Visual-Inertial SLAM |
| Zhang and Singh [4, 5] | 2014/2017 | PC MDC for PC Registration for LO | - | Linear Interpolation | |||||||
| Ceriani et al. [77] | 2015 | LiDAR SLAM | LI | ||||||||
| Dubé et al. [101] | 2016 | LiDAR-Inertial-Wheeled SLAM | Linear Splines | ||||||||
| Droeschel et al. [76] | 2017 | PC MDC for LiDAR SLAM | - | Linear Interpolation | |||||||
| Deschaud [69] | 2018 | PC MDC for LiDAR SLAM | - | Linear Interpolation | |||||||
| Wang et al. [103] | 2018 | Stereo VO | - | Constant-Acceleration Integration | |||||||
| Lowe et al. [104] | 2018 | LiDAR-Visual-Inertial SLAM | Uniform Linear Spline | ||||||||
| Park et al. [83] | 2018 | LiDAR-Inertial SLAM | Linear Interpolation | ||||||||
| [97] | 2018 | Multi-Sensor Odometry and Localization | - | Linear Interpolation | |||||||
| Eckenhoff et al. [98] | 2019 | Multi-Camera VIO and Calibration |
| Ye et al. [71] | 2019 | PC MDC for LO | - |
| Lin and Zhang [70] | 2020 | PC MDC for PC Registration for LO | - | Linear Interpolation | |||||||
| Shan et al. [72] | 2020 | PC MDC for LiDAR-Inertial SLAM | - | Linear Interpolation | |||||||
| Park et al. [86] | 2020 | LiDAR-Visual Calibration | Linear Interpolation | ||||||||
| Park et al. [84] | 2021 | LiDAR-Visual-Inertial SLAM | Linear Interpolation (& Uniform Cubic B-Spline) | ||||||||
| Burnett et al. [80] | 2021 | Spinning RADAR MDC | - | Linear Interpolation | |||||||
| Li et al. [73] | 2021 | PC MDC for LIO | - |
| Wang and Ma [105] | 2021 | Visual-Inertial-LiDAR Extrinsic Calibration |
| Shan et al. [75] | 2021 | PC MDC for LiDAR-Visual-Inertial SLAM | - | Linear Interpolation | |||||||
| Xu and Zhang [78] | 2021 | LIO & PC MDC |
| Xu et al. [106] | 2022 | LIO |
| Wang and Ma [107] | 2022 | PC MDC for LiDAR-Visual-Inertial SLAM | - | Linear Interpolation | |||||||
| Ramezani et al. [85] | 2022 | LiDAR-Inertial SLAM | Linear Interpolation (& Cubic B-Spline) | ||||||||
| Dellenbach et al. [81] | 2022 | PC MDC for PC Registration for LO | - | Linear Interpolation | |||||||
| He et al. [108] | 2023 | LIO |
| Nguyen et al. [74] | 2023 | PC MDC for LiDAR-Inertial SLAM | - | Linear Interpolation | |||||||
| Zheng and Zhu [82] | 2023 | LO | - | Linear Interpolation | |||||||
| Chen et al. [109] | 2023 | PC MDC for LIO | - | Constant-Jerk Integration | |||||||
| Zheng and Zhu [30] | 2024 | LO | - | Linear Interpolation | |||||||
IV-A4 Nonlinear Interpolation
Not all nonlinear interpolation strategies are generalized by the temporal splines or TGPs. In the pose interpolation method for SLAM proposed by Terzakis and Lourakis [110], they regress a quadratic rational function to the previous five or more poses with equality constraints to enforce interpolation. Rational functions take the form of an algebraic fraction where both numerator and denominator are polynomials (i.e., the quotient of quadratic TBFs in this case).
IV-A5 Numerical Integration
[…]
IV-A6 Gaussian Process IMU Measurement Preintegration
[…]
| Authors | Year | Contributions | ||
| Le Gentil et al. [13] | 2018 | UPMs | ||
| Le Gentil et al. [14] | 2020 | GPMs | ||
| Le Gentil et al. [121] | 2020 | Event-Inertial SLAM with GPMs | ||
| Le Gentil and Vidal-Calleja [122] | 2021 |
| Dai et al. [123] | 2022 | Event-Inertial Odometry & Mapping with LPMs | ||
| Le Gentil and Vidal-Calleja [124] | 2023 |
In follow-up work, Le Gentil and Vidal-Calleja [122] propose Unified Gaussian Preintegrated Measurements to extend GPMs to rotations, and Linear Preintegrated Measurements which use LI for computational speed. Dai et al. [123] propose a corner-based event-inertial odometry and mapping system using LPMs for pose inference at the event times. Most recently, Le Gentil and Vidal-Calleja [124] formulate UGPMs for poses, fusing the rotation and acceleration computation steps. They propose initializing UGPMs with LPMs, with an optional kernel hyperparameter tuning step, and have proven effectiveness in a wide variety of multi-sensor estimation problems. These GP-based IMU preintegration methods have been shown to be more accurate than other methods [10, 11], especially during fast motion. Since these methods use a Gaussian kernel, the computational complexity for GP regression is cubic, however this could be addressed by using exactly sparse kernels (Section III-C).
| Authors | Year | Applications | Time-Invariant Estimates | Time-Varying Estimates | |||||||||||||
| State(s) | State(s) | State Representation(s) | |||||||||||||||
| Kim et al. [125, 126]* | 1995 | Torque Computation, Animation | - | Bézier, Hermite & B-Spline | |||||||||||||
| Gortler and Cohen [127]* | 1995 | Theory | - | Polynomial and Wavelet TBFs | |||||||||||||
| Anderson-Sprecher and Lenth [128] | 1996 | Bearing-Based Object Tracking | - | Cubic B-Spline | |||||||||||||
| Qin [52]* | 1998 | Theory | - | B-Spline | |||||||||||||
| Kang and Park [129] | 1999 | Theory | - | Cubic Splines | |||||||||||||
| Crouch et al. [51] | 1999 | Theory | - | Cubic Splines | |||||||||||||
| Jung and Taylor [130] | 2001 | VIO |
| Bibby and Reid [131] | 2010 | Dynamic SLAM | Non-Uniform Cubic Splines | ||||||||||||||
| Hadzagic and Michalska [132] | 2011 | Object Tracking | - | Non-Uniform Cubic B-Spline | |||||||||||||
| Fleps et al. [133] | 2011 | Visual-Inertial Calibration | Cubic B-Splines | ||||||||||||||
| Furgale et al. [134, 20]* | 2012/2015 |
| Furgale et al. [18] | 2013 | Visual-Inertial Calibration |
| Lovegrove et al. [135]* | 2013 |
Cubic B-Spline | |||||||||||||
| Anderson et al. [136, 137] | 2013/2015 | LiDAR SLAM |
| Oth et al. [138] | 2013 | RS Calibration | TBFs / Uniform Cubic B-Spline | ||||||||||||||
| Sheehan et al. [139] | 2013 | LiDAR Localization | - | Cubic Catmull-Rom Spline | |||||||||||||
| Zlot and Bosse [99, 100] | 2013/2014 | LiDAR-Inertial Odometry | Uniform Linear / Cubic B-Spline | ||||||||||||||
| Rehder et al. [140] | 2014 | LiDAR-Visual-Inertial Calibration | TBFs / Quintic B-Splines | ||||||||||||||
| Alismail et al. [141] | 2014 | PC MDC for PC Registration for LO | - | Uniform Linear to Quintic B-Spline | |||||||||||||
| Anderson et al. [48] | 2014 | TE | - | Polynomial and Wavelet TBFs | |||||||||||||
| Patron-Perez et al. [142] | 2015 |
Cubic B-Spline | |||||||||||||
| Kerl et al. [143] | 2015 | RS RGB-D Odometry | - | Uniform Cubic B-Spline | |||||||||||||
| Mueggler et al. [144] | 2015 | Event Camera Localization | - | Uniform Cubic B-Spline | |||||||||||||
| Rehder et al. [145] | 2016 |
| Rehder et al. [146] | 2016 | Multi-IMU Visual-Inertial Calibration |
| Kim et al. [147] | 2016 | RS SLAM | Cubic B-Spline | ||||||||||||||
| Sommer et al. [148]* | 2016 |
B-Spline | ||||||||||||
| Li et al. [149, 150, 151] | 2016-2018 | Bearing-based Tracking | - | Polynomial and Sinusoidal TBFs | |||||||||||||
| Rehder and Siegwart [152] | 2017 | Visual-Inertial Calibration |
| Vandeportaele et al. [153] | 2017 | RS Localization | - | Uniform Cubic B-Spline | |||||||||||||
| Droeschel and Behnke [154] | 2018 | PC MDC, LiDAR SLAM | Cubic B-Spline | ||||||||||||||
| Mueggler et al. [155] | 2018 | Event-Inertial Localization | Uniform Cubic B-Spline | ||||||||||||||
| Ovrén and Forssén [156] | 2018 | RS Visual-Inertial SLAM | Cubic B-Splines | ||||||||||||||
| Ovrén and Forssén [157] | 2019 | RS Visual-Inertial SLAM | Cubic B-Spline(s) | ||||||||||||||
| Sommer et al. [53]* | 2020 |
| Pacholska et al. [158] | 2020 | Certifiable Range-Based Localization | - | Polynomial and Sinusoidal TBFs | |||||||||||||
| Lv et al. [159] | 2020 | LiDAR-Inertial Calibration | Uniform Cubic B-Spline | ||||||||||||||
| Hug and Chli [31] | 2020 | Visual-Inertial SLAM | Cubic B-Spline(s) | ||||||||||||||
| Lv et al. [160] | 2021 |
| Ng et al. [161] | 2021 | Multi-RADAR Inertial Odometry | - |
| Yang et al. [162] | 2021 | Multi-Camera SLAM | Non-Uniform Cubic B-Spline | ||||||||||||||
| Quenzel and Behnke [54] | 2021 | Surfel Registration for LO | - | Non-Uniform Quadratic B-Splines | |||||||||||||
| Huang et al. [163] | 2021 | VO |
Cubic B-Spline | |||||||||||||
| Persson et al. [45]* | 2021 |
Cubic to Quintic B-Spline | |||||||||||||
| Cioffi et al. [19] | 2022 | Visual-Inertial-GNSS SLAM | Cubic to Sextic Uniform B-Splines | ||||||||||||||
| Hug et al. [164] | 2022 | Stereo Visual-Inertial SLAM | Uniform Cubic B-Splines | ||||||||||||||
| Tirado and Civera [55]* | 2022 |
| Mo and Sattar [165] | 2022 | Visual-Inertial SLAAM |
| Huai et al. [166] | 2022 | RS Visual-Inertial Calibration |
Uniform Quintic B-Splines | |||||||||||||
| Wang et al. [167] | 2022 | Event-Based VO | - | B-Spline | |||||||||||||
| Lang et al. [168] | 2022 | Visual-Inertial SLAAM |
| Lv et al. [169] | 2022 | LiDAR-Inertial Calibration |
| Lv et al. [170] | 2023 | LiDAR-Visual-Inertial SLAAM |
| Lang et al. [171] | 2023 | LVIO | - |
| Li et al. [56]* | 2023 | Theory, UWB-Inertial Localization | Uniform Cubic B-Splines | ||||||||||||||
| Li [172] | 2023 | UWB-Inertial Localization | - | Uniform Cubic B-Spline | |||||||||||||
| Jung et al. [173] | 2023 | MDC for Multi-LiDAR Inertial Odometry | - | Uniform Cubic B-Spline | |||||||||||||
| Nguyen et al. [174] | 2024 | LIO | - |
| Lv et al. [175] | 2024 | LO | - | Uniform Cubic B-Splines | |||||||||||||
| Hug et al. [176] | 2024 |
- | Uniform B-/Z-Splines | |||||||||||||
| Li et al. [177] | 2024 | LiDAR-RADAR-inertial Odometry and Calibration |
IV-B Temporal Splines
A chronologically ordered overview of all surveyed methods using temporal splines is given in Table V.
IV-B1 Origins & Early Mathematical Formulations
Splines have received significant attention in robotics as a CT representation of the state. They were first formulated as piecewise-polynomial by Schoenberg in 1946 [178], inspired by ‘lofting’ where thin wooden strips (called splines) were used for aircraft and ship design, including for construction from templates during World War II [179]. While various spline types have been explored in different robotics applications, B-splines, short for basis splines, are most popular due to their smoothness properties. The mathematics of B-splines was pioneered by Schoenberg, de Boor, and others, with the de Boor-Cox recurrence formula [49, 50] a famous early formulation. From this, Qin [52] was the first to derive a general matrix representation for B-splines. Later, splines were presented as a special case of Temporal Basis Functions (TBFs) where the bases are polynomials [20, 134]. As an alternative to polynomials, wavelet basis functions [48, 127] have been proposed in a hierarchical scheme to account for the varying richness of motion. Sinusoidal temporal bases have also been used in conjunction with polynomials for the problem of certifiable range-based localization [158] or object tracking [149, 150, 151].
IV-B2 Generalization to Lie Groups
Importantly for robotics, splines were formulated for orientations [125, 126, 129] and then general Lie groups [51]. Despite this, many early applications retained a vector spline representation for orientation or pose [133, 134, 20, 18, 138, 139, 140, 141, 145, 146] before splines [135, 142, 144, 147] and general Lie group splines of arbitrary order [148] were applied. Transitioning away from the formulation for orientation was necessary because of the presence of singularities, the loss of continuity [125, 126, 135], and because interpolation does not represent a minimum distance in rotation space [135, 153]. For these reasons, the resultant curves under the vector formulation may not realistically approximate motion. Manifold constraints were rarely enforced before the Lie group formulation, although Fleps et al. [133] did so in their SQP approach.
IV-B3 Choice of Pose Representation
For 3D pose estimation using B-splines, a choice exists between the joint () and split ( or ) representations. Ovrén and Forssén [157] scrutinize this choice for robotic applications and advocate for the split representation in most cases. They showed that under the representation, translation, and orientation are coupled in screw motion, where implicitly “the pose is manipulated by an internal force and torque”, an action called a wrench [180], which may be appropriate in some contexts such as for the trajectories of cameras mounted on a robot arm. In contrast, they argue that the split representation implicitly assumes that an external force and torque actuate the pose. Furthermore, they demonstrate that the analytic acceleration of the joint spline is not the kinematic body acceleration in general. Besides the theoretical justifications, they experimentally show that the split representation converges faster and more accurately for their RS Visual-Inertial SLAM system. More recently, Tirado and Civera [55] make a similar argument; however, they advocate for the joint representation as the general case for rigid-body motion. Hug and Chli [31] prefer the split representation, emphasizing the increased computational effort required in the representation. For orientation, Haarbach et al. [46] and Li et al. [56] advocate for quaternions () over rotation matrices (). This can be attributed to their space efficiency, requiring four variables instead of nine, and their relative computational efficiency. For the same reasons, Haarbach et al. [46] advocate for the unit dual quaternion over for B-spline pose interpolation.
IV-B4 Modern Mathematical Advancements
Sommer et al. [53] devise a computationally efficient method for computing Lie group B-spline derivatives, requiring instead of matrix operations for a spline of order . They further derive a recursive strategy to compute analytic Jacobians for the B-spline and its first and second derivatives in time. This was extended to other Lie groups [55, 56] with Tirado and Civera [55] demonstrating significant computational savings over automatic or numerical differentiation. Finally, Johnson et al. [22, Appendix E] provide analytic Jacobians of general Lie group splines up to the second derivative. Hug and Chli [31] contribute analytic Jacobians for the knots, suggesting the possibility that they could be jointly optimized in future work. Recently, Hug et al. [176] use SymForce [181] to obtain analytic B-spline implementations that are much faster than those derived by Sommer et al. [53]. Persson et al. [45] provide a number of novel theoretical contributions: i) constant-velocity extrapolation through new basis functions, ii) a necessary condition on knot spacing for quaternion splines to express a maximum angular velocity, iii) a simple method to derive and compute integrals of spline derivative functions, iv) a theoretical and experimental investigation of different spline regularizers, and v) an approach for how these integral-type costs should be sampled. Spline regularization has been effective in reducing oscillations [20, 18] and enforcing non-holonomic constraints [163]. Additionally, Cioffi et al. [19] demonstrated with batch optimization that B-splines are more accurate than DT for trajectory estimation, especially when estimating for time offsets between sensors.
IV-B5 Early Applications in Robotics
Among the first robotic applications of splines were position estimation in object tracking [128] and Visual-Inertial Odometry (VIO) [130]. In another early work, Bibby and Reid [131] represented the vehicle’s trajectories and dynamic objects in the environment using cubic interpolating splines in a 2D dynamic SLAM application. Hadzagic and Michalska [132] used cubic B-splines for 2D position estimation in ship tracking. Bosse and Zlot [64, 100] used cubic B-splines for trajectory fitting and optimization while keeping LI as their primary CT representation. Later, TBFs were applied to the estimation of robot velocity in a feature-based LiDAR SLAM context [136, 137], using cubic B-splines in the implementation and a Riemann summation of velocity to obtain the pose estimates.
IV-B6 Explicit Motion Distortion Correction (MDC)
As in LI, B-splines were also used to correct for MDC, including in point cloud registration for LO [141], in a LiDAR SLAM system [154] (extending Droeschel et al. [76]), and RSC for RGB-D odometry [143]. The LO and mapping system presented in Quenzel and Behnke [54] uses a temporal B-spline and Gaussian Mixture Models for local-to-map surfel registration. Lv et al. [160] solve the LiDAR MDC problem for their LIO system by first estimating pose and velocity with IMU integration before fitting a uniform B-spline and using it to correct the motion-distorted point cloud. They use the corrected scans in a feature-based LIO optimization where a B-spline again represents the trajectory. Jung et al. [173] use cubic B-spline interpolation for motion and temporal compensation of multiple LiDARs in their LIO KF.
IV-B7 Localization, Odometry & SLAM
Sheehan et al. [139] use a Catmull-Rom spline for LiDAR localization, and the recent work by Li et al. [56] uses B-splines for UWB-inertial localization. B-splines have been used to process high-rate event camera measurements, which are particularly hard to aggregate for DT systems. This includes event [144] and event-inertial localization [155], and event VO [167]. Leveraging prior work [74], Nguyen et al. [174] propose an iterative, parallelized, on-manifold linear solver for real-time surfel-based LIO. Lv et al. [175] use a ‘spot uncertainty model’ to weight point-to-plane residuals in their LO system. Ng et al. [161] apply B-splines to RADAR-inertial odometry, using radial velocity measurements from multiple radars to form ego-velocity residuals. Adopting uniform cubic B-splines for the trajectory, Li et al. [177] jointly optimize LiDAR point-to-plane, RADAR Doppler velocity, and inertial residuals for multi-sensor odometry. B-splines have also been used as the CT formulation for robot trajectories in many SLAM works. This includes RS visual [135, 142, 147], RS visual-inertial [168], visual-inertial [157, 31, 164], multi-camera [162], LiDAR [154], and LiDAR-visual-inertial [170, 171] SLAM.
IV-B8 Sensor Calibration
TBFs and B-splines have been used for sensor calibration, including estimation of the line delay of a RS camera [138], visual-inertial [18] and LiDAR-visual-inertial [140] extrinsic calibration, and comprehensive simultaneous intrinsic and extrinsic calibration for LiDAR-visual-inertial [145], multi-IMU visual-inertial [146], and visual-inertial [152, 45] configurations. More recently, uniform quintic B-splines have been applied to RS visual-inertial calibration [166], and uniform cubic B-splines to targetless LiDAR-inertial calibration in structured environments [159, 169], both covering many intrinsic and extrinsic parameters.
IV-B9 Online Optimization
While many perform batch optimization, some works operate online using sliding-window optimization (a.k.a. fixed-lag smoothing). The first to do so only used linear splines (equivalent to LI) [99, 100, 101, 104], but Zlot and Bosse [100] also did with cubic B-splines. Persson et al. [45] were among the first to use windowed optimization for real-time odometry. Lang et al. [168] performed RS visual-inertial SLAM in a sliding-window optimization scheme with B-splines but could not run in real-time. Lv et al. [170] extended this method to a real-time LiDAR-Visual-Inertial system, developing marginalization strategies for B-spline odometry and SLAM problems. More recent spline-based odometry systems have continued to develop marginalization strategies [171, 175, 177], with Lang et al. [171] to first to do so with non-uniform splines. Importantly, the use of per-point residuals in these works allows MDC to occur simultaneously to state estimation [170, 171, 175, 177]. Quenzel and Behnke [54], Mo and Sattar [165] who incorporate inertial measurements through cubic splines into DSO [182], and Tirado and Civera [55] who employ non-uniform cubic B-splines for object tracking, also employ sliding-window optimization. In contrast to these optimization approaches, Li [172] devises a probabilistic recursive filter that maintains the control points for a single cubic B-spline segment in its state and can thus run in real-time with uncertainty estimation. Hug et al. [176] take a different approach, proposing a Gaussian Belief Propagation (GBP) optimization framework called Hyperion, where factor graph optimization is achieved through efficient message passing between nodes and factors. While usually slower than existing Nonlinear Least Squares (NLLS) solvers such as Ceres [183], GBP computes variable (inverse) uncertainty during optimization instead of as a post-optimization procedure, allowing strategies for targeted updating (e.g., of non-converged nodes), and is inherently scalable to multi-agent systems.
| Year | Applications | Time-Invariant Estimates | Time-Varying Estimates | |||||||||||||||||
| State(s) | State(s) | State Representation(s) | ||||||||||||||||||
| Tong et al. [57] | 2012 |
WNOA TGP | |||||||||||||||
| Tong et al. [47] | 2013 |
WNOA TGP | |||||||||||||||
| Tong et al. [184] | 2014 | Laser-Based Visual SLAM | WNOA TGP | |||||||||||||||||
| Barfoot et al. [34] | 2014 | Range-Based SLAM | WNOA TGP | |||||||||||||||||
| Yan et al. [185, 186] | 2014/2017 | Range-Based SLAM | WNOA TGP | |||||||||||||||||
| Anderson et al. [35] | 2015 | Range-Based SLAM | WNOA TGP | |||||||||||||||||
| Anderson and Barfoot [59] | 2015 | Visual SLAM | WNOA TGP | |||||||||||||||||
| Dong et al. [187] | 2017 | Visual-Inertial-GNSS SLAM |
| Dong et al. [60, 61] | 2017/2018 |
WNOA TGP | |||||||||||||||
| Mukadam et al. [188, 189] | 2017/2019 |
WNOA TGP | |||||||||||||||
| Rana et al. [190] | 2018 | Learning from Demonstration and Motion Planning | - | WNOA/Learnt TGPs | ||||||||||||||||
| Maric et al. [191] | 2018 | Motion Planning | - | WNOA TGP | ||||||||||||||||
| Warren et al. [192] | 2018 | VO for Teach & Repeat | - | WNOA TGP | ||||||||||||||||
| Warren et al. [193] | 2018 | VO for Teach & Repeat | - | WNOA TGP | ||||||||||||||||
| Tang et al. [62] | 2019 | Odometry | - | WNOJ TGP | ||||||||||||||||
| Judd [194], Judd and Gammell [195] | 2019/2020 | Dynamic SLAM | WNOA TGPs | |||||||||||||||||
| Barfoot et al. [196] | 2020 | Wheeled Robot SLAM | WNOA TGP | |||||||||||||||||
| Wong et al. [197] | 2020 | Hyperparameter Estimation for LiDAR Localization | - | WNOA TGP | ||||||||||||||||
| Wong et al. [198] | 2020 |
- | Singer TGP | ||||||||||||||||
| Kapushev et al. [199] | 2021 | SLAM | RFF-Gaussian TGP | |||||||||||||||||
| Judd and Gammell [200, 201] | 2021/2024 | Dynamic SLAM | WNOA / WNOJ TGPs | |||||||||||||||||
| Li et al. [202] | 2021 | LiDAR-Inertial Extrinsic Calibration and Localization |
| Zhang et al. [203] | 2022 | LiDAR-Inertial-GNSS Odometry | - |
WNOJ TGP | |||||||||||||||
| Zhang et al. [204] | 2022 | Inertial-GNSS Odometry | - |
| Wu et al. [205] | 2022 | FMCW LiDAR Odometry | - | WNOA TGP | ||||||||||||||||
| Burnett et al. [206] | 2022 | LiDAR and RADAR Odometry for Teach & Repeat | - | WNOA TGP | ||||||||||||||||
| Liu et al. [207] | 2022 | Event-Based SLAM for VO | WNOA TGP | |||||||||||||||||
| Wang and Gammell [208] | 2023 | Event-Based SLAM for VO | WNOA TGP | |||||||||||||||||
| Yoon et al. [209] | 2023 | FMCW LiDAR-Inertial Odometry | - | WNOA TGP | ||||||||||||||||
| Le Gentil et al. [210] | 2023 | Event MDC Feature Tracking | - | Squared Exponential TGPs | ||||||||||||||||
| Le Gentil and Vidal Calleja [211] | 2023 | IMU-Pose Extrinsic Calibration |
| Zheng and Zhu [212] | 2024 | LIO | - |
| Johnson et al. [22] | 2024 | Visual-Inertial Localization |
| Lisus et al. [213] | 2024 | RADAR Odometry for Teach & Repeat | - | WNOA TGP | ||||||||||||||||
| Burnett et al. [214] | 2024 | LiDAR-Inertial and RADAR-Inertial Odometry | - |
| Barfoot et al. [215] | 2024 | Certifiable Trajectory Estimation | - | WNOA TGP | ||||||||||||||||
| Zhang et al. [216] | 2024 | LiDAR-Inertial-GNSS-Velocimeter Odometry | - |
| Nguyen et al. [63] | 2024 |
WNOJ TGPs | |||||||||||||||||
IV-C Temporal Gaussian Processes (TGPs)
A chronologically ordered overview of all the TGP methods surveyed in this work is presented in Table VI.
IV-C1 Early Works
The use of GPs for CT state estimation was first realized by Tong et al. [57, 47, 217], where they presented Gaussian Process Gauss-Newton (GPGN), a nonlinear batch optimization method, and applied it to 2D range-based localization [57] and SLAM [47]. In these early formulations, the connection with TBFs was made explicit in the derivation, with the transition from parametric to non-parametric achieved through the kernel trick. Tong et al. [184] subsequently extended the GPGN method to 3D for their laser-based visual SLAM system. GPGN preceded the works of Barfoot et al. [34] and Anderson et al. [35], who demonstrated that GPs using the LTV-SDE class of prior have exactly sparse inverse kernel matrices, enabling efficient regression and interpolation. The computation of the inverse kernel matrix is reduced from cubic to linear time, and interpolation can be computed in constant time. When using this prior with a linear measurement model, this approach is shown to be exactly equivalent to discrete-time smoothing when evaluated at the measurement times with identical computational complexity [35, 24]. Solving the SLAM problem using this continuous-time method was called Simultaneous Trajectory Estimation And Mapping (STEAM), with Anderson and Barfoot [59] extending the approach to poses using local LTV-SDEs (Section III-C1). Nonlinear models are supported in these works through linearization and iterative optimization.
IV-C2 Optimization
Exploiting the sparsity of the formulation, the STEAM problem can be modeled as a factor graph, with binary GP prior factors connecting adjacent states [35, 59]. Yan et al. [185, 186] used this to develop a GP version of the incremental optimizer iSAM2 [218], and with iGPMP2 [58] applied this incremental GP approach to motion-planning problems. Sparse interpolated measurement factors allow measurements to be added between the explicitly estimated states [47, 35, 186, 36], and can be readily used within an incremental Bayes tree optimization scheme [186]. Dong et al. [60, 61] provided a formulation for general matrix Lie groups before integrating it into their Simultaneous Trajectory Estimation And Planning (STEAP) framework [188, 189].
Exploring a different direction, Barfoot et al. [196] propose Exactly Sparse Gaussian Variational Inference (ESGVI) to fit a Gaussian distribution to the full Bayesian posterior of the problem states. This approximation of the posterior ought to be more representative than the Laplace approximation obtained from MAP estimation (Equation 6). They devise a Newton-style iterative solver that avoids the need for analytic derivatives through cubature rules and exploit the inherent sparsity of the problem structure to achieve the same computational complexity as MAP estimation. Wong et al. [197] then extend this approach for hyperparameter learning with an Expectation Maximization (EM) optimization framework.
IV-C3 Applications using the WNOA Prior
The GN TGP approach [35, 59, 60] has been applied in the context of teach and repeat, such as in Visual Odometry (VO) for UAVs [192, 193], or LiDAR and RADAR odometry [206, 213]. Liu et al. [207] and Wang and Gammell [208] have proposed incremental event-based VO pipelines based on this CT representation, both employing the GN algorithm in the underlying optimization. Li et al. [202] demonstrated LiDAR-inertial extrinsic calibration and localization in structured environments. Incremental optimizers based on the Bayes tree [218] have been used in SLAM for agricultural monitoring [187] and inertial-GNSS odometry [204]. Two recent works [205, 209] have applied TGPs to odometry with Frequency-Modulated Continuous-Wave (FMCW) LiDARs. Rather than aggregating LiDAR points for registration or feature extraction, Wu et al. [205] formulate point-to-plane and radial-velocity residuals for each point. By contrast, Yoon et al. [209] combine the radial-velocity residuals with a gyroscope for linear estimation of the total DoF ego-velocity, which they integrate into poses. Judd and Gammell [195, 194] address the multimotion estimation problem with TGPs in their occlusion-aware VO pipeline, estimating for the ego-motion and trajectories of points on dynamic objects.
IV-C4 Applications using Other Priors
The aforementioned works all use the WNOA prior. Tang et al. [62] propose the ‘constant-acceleration’ White-Noise-On-Jerk (WNOJ) model instead. They achieve higher accuracies in some cases and prove that when the chosen prior does not match the motion, the state estimate will be biased along certain DoF. The WNOJ prior has since been used for online GNSS multi-sensor odometry [203, 216]. Recently, Zheng and Zhu [212] propose multi-sensor LIO using WNOJ for position, WNOA for orientation, and White-Noise-On-Velocity (WNOV) (a.k.a. ‘random-walk’) for IMU biases, and achieve state-of-the-art results. As in Wu et al. [205] and their previous work [30], they avoid measurement aggregation (e.g., LiDAR registration, IMU pre-integration) while retaining real-time performance through the split pose representation and efficient map management. Burnett et al. [214] also utilize the WNOV prior for IMU biases, demonstrating effective LiDAR-inertial and RADAR-inertial odometry on various datasets. Judd and Gammell [200, 201] compare the WNOJ and WNOA priors in their Multimotion Visual Odometry (MVO) pipeline [194, 195]. Johnson et al. [22] recently were the first to compare TGPs and B-splines in visual-inertial localization experiments. They compared WNOJ TGPs to cubic and quintic B-splines, both with and without motion prior factors, and for both the split and joint pose representations. They conclude that when the same motion models and measurements are used and the degree-of-differentiability matches, temporal splines and GPs are similar in accuracy and solve time. Furthermore, they show that motion prior factors act as regularizers, preventing overfitting to noisy measurements. However, the benefit of these is significantly diminished when IMU is available. Zheng and Zhu [30] and Johnson et al. [22] investigate the efficacy of the split pose representation and conclude its primary benefit is the reduction of computation time. Most recently, Nguyen et al. [63] open-sourced their split pose WNOJ implementation and provide localization and calibration examples for a variety of multi-sensor combinations. They further provide a generalization of White-Noise-On-Derivative (WNOD) priors for any order and analytic Jacobians for all their residuals. Finally, Wong et al. [198] investigate the Singer prior, which represents latent accelerations as a GP with a Matérn kernel. They argue that the data-driven optimization of an additional parameter enables a better representation of the robot’s motion.
IV-C5 Other Approaches
Le Gentil et al. [210] propose a novel method for event camera motion compensation using GPs. By treating camera motion as motion in the image frame over short time periods, they model each DoFs as independent zero-mean GPs with Gaussian kernels. The inducing values (i.e., set of observations) of these are then the hyperparameters of another GP representing a continuous occupancy field of events in the image plane. By optimizing these through log-marginal likelihood maximization, this occupancy field, and hence a continuous distance field as its negative logarithm, can be inferred and used for homography registration and pattern tracking. Le Gentil and Vidal Calleja [211] use a GP pose representation in the extrinsic calibration of an IMU with pose measurement reference frame (e.g., motion capture or robotic arm frames). Kapushev et al. [199] propose a SLAM algorithm evaluated in 2D, which models the pose as a GP with a Random Fourier Features (RFF) approximation of the Gaussian kernel and uses either a discretized motion model or smoothing splines for the prior mean function.
V Open Problems
V-A Temporal Splines
V-A1 Knot Selection
The optimal selection of knots remains one of the most important unresolved problems of the spline method. The simplest and most common approach is to use uniform splines, tuning the frequency according to the application. While uniform splines benefit from a fixed blending matrix that needs to be computed only once, they implicitly assume that excitation of the state evolution (e.g., robot motion) remains constant. Hence, uniform knots can lead to underfitting or overfitting of the trajectory in settings with varying dynamics [31]. Furthermore, a control point is only well-constrained if there are sufficient measurements on the associated segments, meaning that optimization can be poorly posed if the knot frequency is too high relative to the measurement frequency.
Some works associate knots with the measurement [131, 162], with Bibby and Reid [131] selectively removing them where acceleration and jerk are sufficiently similar. Others [153, 171] increase the knot frequency according to the maximum magnitudes of acceleration and angular velocity (from IMU measurements). Another suggestion is to iteratively add knots until the mean visual reprojection error is sufficiently low [153] or to increase knots until the reprojection residuals within segments agree with an expected value [138]. Anderson et al. [48] use this idea to decide when to increase the active resolution level in their hierarchical formulation, based on work by Gortler and Cohen [127] who apply wavelet detail functions over a base B-spline signal function.
An alternative approach is spline error weighting [156, 157], which recognizes that existing methods neglect approximation error arising from model mismatch. By formulating spline fitting in the frequency domain, they devise a method to capture model mismatch and measurement error by setting the residual covariance based on the measurement noise and variance of the error. They further use this frequency analysis to propose a method for selecting a suitable uniform knot frequency based on the fraction of signal energy to be retained.
Several alternative schemes have been offered. Dubé et al. [101] propose several distributions for the knots within the sliding window of the optimization. These include an exponential distribution, a frequency analysis distribution based on average correction signal power, and a uniform distribution, potentially in combination. Hug and Chli [31] discuss optimization of the knots themselves and derive Jacobians for them. However, this idea has not yet been pursued experimentally. Regularizing motion terms [20, 45, 22] may avoid the issue of under-constrained control points; however, it may add assumptions conflicting with the implicit spline constraints.
V-A2 Certifiability & Recovery Guarantees
Important open research questions arise because the optimization methods commonly used for estimation may converge to bad local minima or fail to converge altogether. Indeed, depending on the measurement modalities used, the residuals of the factor graph (i.e., optimization cost terms) may be non-convex. Particularly, if states are confined to manifold spaces, this may lead to non-convex constraints and exacerbate the problem. Multiple recent works [219, 220, 221] show that non-convexity may give rise to bad local minima in which local solvers such as GN can get stuck. Estimating splines may also suffer from this, although the inherent regularizing properties could help escape local minima.
A related question is that of the uniqueness of the solution. Because splines promote smoothness, fewer measurements may be required to obtain unique state estimates. However, exactly how many are necessary remains an open problem, and failure to meet this may result in ambiguities. Recovery guarantees have been derived for polynomial or sinusoidal basis functions with noiseless range measurements [158], but extending these to more general cases remains open.
V-A3 Incremental Optimization
Sliding-window optimization has been used to facilitate real-time operation [99, 100, 101, 104, 54, 165, 55, 168, 170, 171, 174, 175]. Other incremental schemes, such as those based on the Bayes tree [218], have not been explored. Various graphical models have been presented for splines [164, 165, 55], however only very recent works [168, 170, 171, 175] have made serious efforts to represent spline estimation with factor graphs and use these to design marginalization strategies. Leveraging these to incorporate incremental optimizers like iSAM2 [218] could be explored in future work.
V-A4 Initialization
Initialization of control points is rarely discussed in the literature. However, appropriate initialization certainly affects convergence time and optimality. Simple strategies for initializing new control points include using the value of the previous control point, or IMU dead reckoning. However, the state estimate may be far from the control point, especially for highly dynamic motion and high-order splines, and no inverse interpolation function exists to compute control points from the state. One strategy adopted by Huang et al. [163] was to use a spline approximation algorithm from Piegl and Tiller [222] for initialization. Li et al. [177] devise a sensor-specific coarse-to-fine procedure to initialize control points, biases, and extrinsic calibration parameters. Kang and Park [129] offer initialization equations for different parameterizations of orientation splines based on boundary conditions. Yet, initialization remains an open problem.
V-A5 Covariance Interpolation
It is often desirable to infer the uncertainty of the state at the interpolation time from the uncertainty of the control points. In fact, the covariance can be interpolated for vector-valued states, with some limited empirical results demonstrated [20, Figure 6]. Bibby and Reid [131] provide a derivation for their non-uniform cubic splines, which they applied to dynamic SLAM in 2D, using the Jacobian of the spline interpolation function. State covariance is also estimated by the probabilistic filter proposed by Li [172]. However, no works have provided a covariance interpolation formulation for Lie group splines.
V-B Temporal Gaussian Processes (TGPs)
V-B1 State Time Selection
The TGP method provides flexibility in selecting state times. This is analogous to the spline knot selection problem. However, several differences in the methodology must be considered. First, the process state variables are always constrained through GP prior factors, meaning that TGPs are less sensitive to sensor dropout and do not overfit like splines. Additionally, while measurements can be incorporated at any time, there are computational advantages of using non-interpolated factors, achieved by setting the state time to the measurement time. Finally, when using local LTV-SDEs, the approximation error of the linearization is more significant when the changes between states are more considerable. Thus, intuitively, the temporal density of states should be higher when the motion is more dynamic and less when the state changes slowly. Regardless, the approaches proposed for spline knot selection may also apply to TGPs.
A straightforward approach is to set the state times as the measurement times of low-rate sensors, exploiting the efficiency of non-interpolated factors. This is more advantageous when many factors are associated with a single measurement time, such as landmark measurements in global shutter visual SLAM. Adapting the state times based on IMU measurements or residual errors as a surrogate metric for motion change would also be possible. For more theoretical approaches, increasing knots until residual errors match expected values [138] could be applied to state times. Alternatively, the state times - usually fixed dependent variables - could be added as optimization variables, as discussed for knots [31]. Furthermore, residuals could be weighted analogously to spline error weighting [156, 157] to account for the model approximation error uncaptured by the LTV-SDE prior.
V-B2 Certifiability & Recovery Guarantees
As for splines, common measurement models and state constraints may yield non-convex optimization problems that challenge convergence and global optimality guarantees for local solvers such as GN. A series of works on (DT) globally optimal state estimation [223, 221, 220] shows that many MAP problems can be formulated as Quadratically Constrainted Quadratic Programs, allowing for tight (convex) Semidefinite Program (SDP) relaxations. The latter can be used to globally solve or certify CT estimation problems.
In contrast with the spline method, GP motion priors can be effortlessly integrated with this paradigm, as shown recently for the examples of CT range-only localization [224] and pose-graph optimization [215]. Indeed, to incorporate a GP motion prior, the state is augmented with the necessary number of temporal derivatives, after which the motion prior regularization always takes the form of additional inter-state quadratic cost terms, which can be readily included in the QCQP formulation. However, augmenting the state and regularizing the problem in this way may affect the tightness of the SDP relaxations of the QCQP, which, in turn, can break certifiability. For pose-graph optimization in particular, a significant number of so-called redundant constraints are required for tightness at reasonable noise levels when using the WNOA prior [215, Equation (57)], which breaks typical constraint qualifications and significantly increases the computational complexity of the used SDP solvers. Measures to increase and (a-priori) understand tightness and to speed up SDP solvers are crucial for advancing the practicality of these methods.
As for the required number of measurements, TGPs behave differently than splines because of the Bayesian nature of the approach. For example, in the absence of measurements, the method will return a trajectory according to the motion prior. This allows the method to interpolate between states even when no measurements are acquired, while the uncertainty of such interpolations will grow.
V-B3 Complex Process Models
Tang et al. [62] show that selecting a process model prior that does not represent state evolution leads to biased estimation. Thus, it is clear that a prior that most appropriately reflects the state evolution should be chosen. The TGP method inherently allows for complex process dynamics and control inputs to regularize the trajectory and guide interpolation. However, thus far, only relatively simple models, such as the WNOA or WNOJ priors, have been examined. While they can offer computational savings and require tuning only a small number of hyperparameters, they may not accurately reflect the process dynamics of complex robotics systems or even simple processes from other fields. Questions about how beneficial and computationally tractable more complex priors would be and how feasible these would be for real-time applications, remain open for this method.
Some recent works have used learning in the modeling of system dynamics for fast online use [225, 226], and it would be reasonable to consider this idea in the computation or augmentation of the prior. A principled method for tuning hyperparameters (e.g., ) is the maximization of the marginal log-likelihood using a training data set [35] (over hand-tuning [62, 58, 216] or fitting to data [59]), such as with the Singer prior [198]. While this typically requires ground truth data, the EM parameter learning framework extending ESGVI [196] proposed by Wong et al. [197] demonstrates that hyperparameters can be learned from the original noisy measurement data. The Singer prior [198] is an example of an alternative approach to process modelling, whereby an SDE is derived from factorization of the spectral density of a GP prior, such as a kernel from the Matérn family [227, 228]. It remains to be seen what methods could be effectively employed to learn reasonable GP priors for complex processes.
VI Applications to other Domains
VI-A Temporal Splines
Splines have been extensively used for planning in robotics, including wheeled robots [229], legged locomotion [230], wheel-legged locomotion [231], and manipulation [232, 233]. However, due to their smoothness guarantees, splines are especially appealing for differentially flat systems, such as Unmanned Aerial Vehicles [234, 235]. In a differentially flat system, the states and the inputs can be written as algebraic functions of some variables (called flat outputs) and their derivatives. The continuity of splines can guarantee that the UAV’s trajectory is dynamically feasible.
Moreover, splines whose basis functions form a partition of unity (i.e., they are nonnegative, and their sum is one) exhibit the convex hull property, which guarantees that each spline segment is wholly contained within the convex hull of its control points. This is leveraged by many trajectory planning works [236, 237, 238, 239, 240, 241, 242], since it substantially simplifies obstacle avoidance constraints [243]. B-splines are the typical choice of splines for trajectory planning, as they can guarantee smoothness and high continuity by construction without imposing explicit constraints. However, the convex hull of the B-Spline control points is typically much larger than the segment itself. To reduce conservativeness, some works [244] use B-spline control points as decision variables in trajectory planning. However, the Bézier control points of each interval of this B-spline are used in the obstacle avoidance constraints. The MINVO control points [245, 246] allow for even tighter enclosures and have been leveraged for obstacle avoidance [246].
To a limited degree, shape estimation for continuum or flexible robots has also been found to be amenable to spline and TBF theory since they are also one-dimensional estimation problems, parameterized by arc length instead of time. For example, weighted combinations of basis functions [247, 248] and Bézier curves [249] have been explored.
Splines have also been leveraged to represent specific environment features. For instance, Catmull-Rom splines are used by Qiao et al. [250] to parameterize the road lanes. B-splines are used by Rodrigues et al. [251] to represent unstructured 2D environments more compactly, and in other works [252, 253, 254] to represent the sensor data in their state estimation frameworks. Spline theory has also been utilized for perception outside the state estimation problems discussed in this work, such as the spatio-temporal representation of event camera feature tracks with Bézier curves [255].
In machine learning, apart from traditional regression [256, 257], splines have also recently found applications in neural networks [258, 259, 260], including recently Kolmogorov-Arnold Networks [261], proposed as a promising alternative to the traditional Multilayer Perception (MLP). Moreover, Yang et al. [262] and Roth et al. [263] use splines as the output of their end-to-end learned neural network-based local planner, trained imperatively through Bilevel optimization.
VI-B Temporal Gaussian Processes (TGPs)
GPs have been deployed for a large number of other tasks in robotics. Zhang and Scaramuzza [264] demonstrates the capacity for temporal GPs with Gaussian kernels to represent ground truth trajectories, proposing generalized absolute and relative errors for principled trajectory evaluation. GPs with Gaussian kernels have also been used in Bayesian filters, such as for replacing or enhancing the parametric prediction and observation models [265, 44, 266]. The use of GPs as surrogate observation models has been explored for signal-strength (e.g., WiFi, magnetic flux) state estimation applications [267, 268, 269, 270], and as dynamics prediction models within the control community [271, 272, 273, 274, 275]. In informative path planning applications, GPs have been used for representing scalar environmental fields, like temperature, salinity, methane concentration, magnetic field intensity, and terrain height [276, 277, 278, 279, 280, 281]. GPs have also seen extensive use in mapping, including occupancy maps [282, 283, 284, 285, 286, 287, 288, 289, 290, 291, 292, 293], implicit surfaces [294, 295, 296, 297, 298, 299, 300, 301, 302, 303, 304, 305], their combination [306], gradient maps [307, 308, 309, 310] and distance fields [302, 305, 311, 210, 312, 313]. They have also been used as a representation for local perception and exploration uncertainty [314]. However, these works did not use LTV-SDE priors to obtain exact sparsity in GP regression.777Although, some works [306] attain sparsity in the covariance matrix by using a piecewise kernel that is zero beyond a certain distance [315].
Mukadam et al. [316] were the first to apply exactly sparse TGPs to the problem of motion planning with the GPMP algorithm, in particular for a 7-DoF robotic arm, addressing the large number of states needed by discrete-time planners. This was extended to factor graph optimization, in GPMP2 [317] (later used by Maric et al. [191]), GPMP-GRAPH [318], and iGPMP2 [317, 58], before unification with state estimation in STEAP [188, 189], and learning from demonstration in CLAMP [190]. CLAMP is unique from other works in that the transition matrix and bias are learned for particular tasks from trajectory demonstrations with linear ridge regression. GPMP2 was also extended to dGPMP2 [319], which through a self-supervised end-to-end training framework, allowed the learning of factor covariance parameters, such as , an idea which could be applied to state estimation in future work. Inspired by these works, Cheng et al. [320] adopts an exactly sparse GP for path planning in an autonomous driving context, parameterizing their WNOJ prior by arc length instead of time. Lilge et al. [321] similarly parameterizes by arc length to model the shape of a continuum robot using a white-noise-on-strain-rate (analogous to WNOA) LTV-SDE prior derived from the Cosserat rod model, with state consisting of pose and generalized strain.
VII Conclusion
This work examined the literature and methods of continuous-time state estimation for robotics. Interpolation and integration techniques remain popular CT methods, especially linear interpolation due to its simplicity and low computation cost. This work has focused more heavily on temporal splines and Gaussian processes, which can represent complex process dynamics while remaining computationally competitive. By modeling state evolution explicitly, rather than capturing ‘snapshots’ at measurement times, CT methods offer significant advantages over DT estimation, both for the performance, flexibility, and generalizability of the estimator itself and for downstream planning or control algorithms. By providing introductory theory, a comprehensive survey, and a discussion of open problems and applications in other domains, this work aims to support future CT research, which will become increasingly essential to the field as the number, variety, and complexity of sensors grow in robotics applications.
| Variable | State Representation(s) | Description |
| -Dimensional Position | ||
| -Dimensional Linear Velocity | ||
| -Dimensional Linear Acceleration | ||
| -Dimensional Orientation | ||
| -Dimensional Angular Velocity | ||
| -Dimensional Angular Acceleration | ||
| -Dimensional Pose | ||
| -Dimensional Pose Velocity | ||
| -Dimensional Pose Acceleration | ||
| - | Sensor Calibration Variables | |
| - | Sensor Intrinsics | |
| - | Camera Projection Intrinsics | |
| Rolling Shutter Line Delay | ||
| Camera Exposure Time | ||
| Camera Response Parameters | ||
| Camera Distortion Parameters | ||
| Range Bias | ||
| Range Scale | ||
| - | Range Frame Misalignment Parameters | |
| Gyroscope Bias | ||
| Gyroscope Scale | ||
| Gyroscope Axis Misalignment | ||
| Gyroscope G-Sensitivity | ||
| Accelerometer Bias | ||
| Accelerometer Scale | ||
| Accelerometer Axis Misalignment | ||
| GNSS Receiver Clock Bias | ||
| GNSS Receiver Velocity Bias | ||
| Number of Satellites | ||
| - | Sensor Extrinsics | |
| Inter-sensor Relative Poses | ||
| IMU-Accelerometer Translation | ||
| Gyroscope-Accelerometer Rotation | ||
| Inter-sensor Time Offsets | ||
| Gravity Vector | ||
| Gravity Orientation | ||
| Map Scale | ||
| - | Map in -Dimensional Space | |
| - | Surfel Map in -Dimensional Space | |
| - | Occupancy Map in -Dimensional Space | |
| Point Landmarks in -Dimensional Space | ||
| Plane Landmarks in 3D Space | ||
| Pixel/Landmark Depths/Distances | ||
| Pixel/Landmark Inverse Depths/Distances | ||
| Pixel/Landmark Bearing Vectors | ||
| Material Optical Parameters | ||
| / | / | Dynamic Object Poses/Velocities |